A passengers train's speed is 60 mi/h, and a freight trains speed is 40 mi/h. The passenger train travels the same distance in 1.5 hours less time than the freight train. How long does each train take to make the trip?
Answer by richwmiller(17219)

60(x-1.5)=40x
60x-90=40x
20x=90
x=4.5 hrs freight
x-1.5=3 hrs. passenger
check
60*3=40*4.5
180=180
ok